OpenStudy (anonymous):

Solve 5x − 6y = −38 3x + 4y = 0 (4, 3) (−4, 3) (4, −3) (−4, −3)

OpenStudy (anonymous):

5x - 6y = -38 5x = 6y - 38 x = 6/5y - 38/5 3x +4y = 0 4y = -3(6/5y - 38/5) 4y = -18/5y + 114/5 4y + 18/5y = 114/5 20/5y + 18/5y = 114/5 38/5y = 114/5 Multiply both sides by 5 to get ride of the denominator 38y = 114y 38/38y = 114/38 y = 3 x = 6/5y - 38/5 x = 6/5(3) - 38/5 x = 18/5 - 38/5 x = -20/5 x = -4 (-4,3) Answer is B, there's the work.

OpenStudy (anonymous):

ty